I think of a number A.
When A is divided by 3, I get B.
When 40 is added to B, I get C.
When C is multiplied by 2, I am left with a number D, which is less than 100.
What is A?

Answer by swincher4391(1107)

A/3 = B
B + 40 = C -> A/3 + 40 = C
C * 2 = D (which is less than 100) -> (A/3 + 40)*2 = D (which is less than 100)
2/3 A + 80 < 100
2/3 * A < 20
A < 30
Assuming that A,B,C,and D are all integers, A can be any multiple of 3 less than 27. Even negative integers theoretically work. There is not a unique solution but any number following that form will work. If being an integer is not a requirement (you may want to check), then the only requisite is that A<30. Even 29.9999 works.
Tests theory:
Say A = 6.
A=6
B=2
C=42
D=84<100
Say A = -15
A = -15
B = -5
C = 35
D = 70 < 100
etc.